Buffer solution was created containing 1.00M acetic acid and 0.500M sodium acetate. The Ka of acetic acid is 1.77x10^-5. What is the pH of the solution
1
Expert's answer
2020-06-11T10:27:20-0400

pH of a solution is calculated using the formula below,

"pH=pKa+log\dfrac{[base]}{[acid]}"


"Ka=1.77 * 10^{-5}"

lets calculate pKa

"pKa=-logKa"

"=-log1.77*10^{-5}"

"=4.752"

therefore,

"pH=4.752+log\dfrac{[0.50]}{[1.00]}"


"=4.752-0.301"


"pH=4.451"